# Improving the mean-field approximation in continuous models of population dynamics with nonlocal dispersal: applications to vegetation pattern formation

### v1

@@ 3.2 Pattern formation in the field equation
We can perform a linear stability analysis of this equilibrium for small spatiotemporal perturbations, considering that $Z_{2}$ (and hence $f$) is given. This calculation provides the linear growth rate of the density Fourier modes
@@
where $F[]$ indicates a Fourier transform operation. The onset of pattern formation is defined by the parameter values such that $\lambda(\mathbf{k}_{c})=0$, where $\mathbf{k}_{c}$ is the non-zero Fourier mode at which the perturbation growth rate is maximum. Therefore, we can define the region where patterns might exist by scanning the $s,r_{c}$ parameter space and testing for the pattern onset condition.
@@
To quantify how accurately this linear stability analysis predicts the region of the IBM in which regular patterns form, we need to measure regularity in the stochastic patterns obtained from the discrete simulations. We introduce two different Fourier-based metrics, a first one based on comparing the distribution of Fourier spectra generated by simulated and random patterns, and a second one based on how strongly the pattern wavenumber (maximum of the pattern power spectrum) contributes to the total power of the pattern spectrum (see Appendix A for details on how to compute these metrics). As expected, both metrics identify that regular patterns form when the dispersal range is short and the competition range is large (bottom right corner in Fig. 4). When comparing these results with the predictions of the field equation, we find that the mean-field approximation (gray line in Fig. 4) underestimates the region of the $s,r_{c}$ parameter space where periodic patterns exist. However, accounting for spatial correlations, either using the SMD equations or IBM simulations, provides a much better agreement with the results of the discrete model (cyan symbols and red dashed line in Fig. 4). Additionally, the agreement between the corrected PDE and the IBM simulations is much better when regularity in the discrete patterns is measured using the relative amplitude of the pattern wavenumber. This metric provides a much more robust measurement of regularity because it does not rely on any statistical modeling assumption. The first metric we introduced, however, quantifies regularity as a statistically significant difference between random patterns and those resulting from the IBM simulation, which makes the definition of regularity sensitive to both the number of pattern realizations we used to construct the statistical ensembles and the percentiles used to define the difference between power spectra distributions. As a result, the transition from regular to non-regular patterns, indicated by the white region in Fig. 4a, will vary depending on how we choose these quantities.
Figure 4: Validation of how the different PDE approximations predict regular patterns. A) Regularity is quantified by comparing IBM simulated and random patterns. B) Regularity is measured by the power of the pattern wavelength normalized by the total power (see App. A for details on each metric). Curves in the panels correspond to each approximation in the treatment of the spatial correlations (see legend in panel (B).
@@ 4 Conclusions
We investigated the spatial patterns that form in a spatially explicit IBM logistic model with nonlocal dispersal and intraspecific competition using both a Gaussian and top-hat competition kernel. Combining numerical simulations of the discrete model and a spatial moment dynamics approximation, we found that, for top-hat competition kernels, the population might exhibit periodic patterns with aggregates of individuals forming a hexagonal lattice [46]. This type of pattern is not present when long-range competition is modeled using a Gaussian kernel [36]. Interestingly, the emergence of these regular patterns also affects the system productivity, allowing for larger population sizes even in conditions where local individual crowding is high [51].
@@ Acknowledgments
Appendices
App. A Computation of the regularity metrics
We introduced two different Fourier-based metrics to quantify regularity in the IBM spatial patterns:
[…]
  • The second metric quantifies the periodicity in the IBM simulated patterns measuring the amplitude of the characteristic wavenumber (the wavenumber at which the radial power spectrum is maximum) to the total power, computed as the integral of the radial power spectrum across the entire range of wavenumber.
[…]

### v2

@@ 3.2 Pattern formation in the field equation
We can perform a linear stability analysis of this equilibrium for small spatiotemporal perturbations, considering that $Z_{2}$ (and hence $f$) is given. We insert the expressions $\langle n(\mathbf{x})n(\mathbf{x^{\prime}})\rangle=\langle n(\mathbf{x})\rangle\langle n(\mathbf{x^{\prime}})\rangle\left(1+f(\mathbf{x}-\mathbf{x^{\prime}})\right)$ and $\langle n(\mathbf{x})\rangle=\langle n\rangle_{h}+\Delta n\,e^{i\mathbf{k}\cdot\mathbf{x}+\lambda(\mathbf{k})t}$ in Eq. (2.21), where $\Delta n$ is the amplitude of the spatiotemporal perturbation. Then, linearizing in $\Delta n$, we obtain an equation describing the linear dynamics around the equilibrium. For more details of this technique applied to ecosystem models, see Ref. [65]. This calculation provides the linear growth rate of the density Fourier modes
@@
where $F[]$ indicates a Fourier transform operation. The onset of pattern formation is defined by the parameter values such that $\lambda(\mathbf{k}_{c})=0$, where $\mathbf{k}_{c}$ is the non-zero Fourier mode at which the perturbation growth rate is maximum. Therefore, we can define the region where patterns might exist by scanning the ($s,r_{c}$) parameter space and testing for the pattern onset condition.
To quantify how accurately this linear stability analysis predicts the region of the IBM in which regular patterns form, we need to measure regularity in the stochastic patterns obtained from the discrete simulations. We introduce two different Fourier-based metrics, a first one based on comparing the distribution of Fourier spectra generated by simulated and random patterns, and a second one based on how strongly the pattern wavenumber (maximum of the pattern power spectrum) contributes to the total power of the pattern spectrum.
[…]
The second metric quantifies how strong is the periodicity in the IBM simulated patterns measuring the amplitude of the characteristic wavenumber (the wavenumber at which the radial power spectrum is maximum) compared to the total power, computed as the integral of the radial power spectrum across the entire range of wavenumber. The reason is that, a perfectly linear periodic pattern (a sine function, for example) has all the power in the peak of the spectrum, and a distribution with no periodicity shows a flat power spectrum.
[…]
@@
Both metrics identify that regular patterns form when the dispersal range is short and the competition range is large (bottom right corner in Fig. 5). To compare these results with the predictions of density equations, both mean-field and accounting for spatial correlations, we evaluated the pattern formation onset derived from Eq. (3.2) for the different approximations of $f(\mathbf{r})$ quantifying the density correlations. To perform this calculation, we imposed the Turing instability condition, $\lambda(\mathbf{k_{c}})=0$ and $\nabla\lambda(\mathbf{k_{c}})=0$, in Eq. (3.2) for the three different approximations of the $f(\mathbf{r})$ function: $f=0$, which corresponds to the mean-field approximation neglecting spatial correlations, $f(\mathbf{r})$ obtained by evaluating the spatial covariances in Eq. (2.30) using IBM simulations, and $f(\mathbf{r})$ obtained from the SMD approximation as defined by Eq. (2.31). The mean-field approximation (gray line in Fig. 5) underestimates the region of the $(s,r_{c})$ parameter space where periodic patterns exist. However, accounting for spatial correlations, either using the SMD equations or IBM simulations, provides a much better agreement with the results of the discrete model (cyan symbols and red dashed line in Fig. 5). Additionally, the agreement between the corrected PDE and the IBM simulations is much better when regularity in the discrete patterns is measured using the relative amplitude of the pattern wavenumber (second metric). This metric provides a much more robust measurement of regularity because it does not rely on any statistical modeling assumption. The first metric we introduced $\mathcal{R}$, however, quantifies regularity as a statistically significant difference between random patterns and those resulting from the IBM simulation, which makes the definition of regularity sensitive to both the number of pattern realizations we used to construct the statistical ensembles and the percentiles used to define the difference between power spectra distributions. As a result, the transition from regular to non-regular patterns, indicated by the white region in Fig. 5a, will vary depending on how we choose these quantities.
Figure 5: Validation of how the different PDE approximations predict regular patterns. A) Regularity is quantified by measuring statistically significant differences between IBM simulated and random patterns (see main text for a detailed description of how the metric is computed). B) Regularity is measured by the relative amplitude of the pattern wavenumber (i.e., ratio between the power of the wavenumber with maximum power and the total power). Curves in both panels show the onset of pattern formation obtained from the linear stability analysis of the density equation. Each curve corresponds to a different approximation to treat the spatial correlations as indicated in the legend in panel (B).
@@ 4 Conclusions
We investigated the spatial patterns that form in a spatially explicit IBM logistic model with nonlocal dispersal and intraspecific competition using both a Gaussian and top-hat competition kernel. Combining numerical simulations of the discrete model and a spatial moment dynamics approximation, we found that, for top-hat competition kernels, the population might exhibit periodic patterns with aggregates of individuals forming a hexagonal lattice [46]. This type of pattern is not present when long-range competition is modeled using a Gaussian kernel [36], this fact is due to our IBM rules, and can be seen mathematically by inserting a Gaussian kernel in Eq. (3.2) for $f=0$ (mean-field approximation); however, there are models for which a Gaussian competition kernel can produce patterns, highlighting that the issue of the general conditions for pattern formation is still debated [33]. Interestingly, the emergence of these regular patterns also affects the system productivity, allowing for larger population sizes even in conditions where local individual crowding is high [52].
@@
For ecological modeling applications in general and particularly in the case of vegetation patterns, individual-based models provide a robust framework with a high degree of accuracy in simulating individual-level processes and interactions, capturing dynamics down to the resolution of single plants. Consequently, these models are very effective for understanding the fine-scale mechanisms driving vegetation pattern formation. However, the computational overhead in simulating these IBMs becomes prohibitively expensive when the system size is large or longer temporal scales are considered, limiting their practical applicability. Density-based models, on the other hand, are computationally very efficient and often amenable to analytical exploration. However, these models assume that interactions are, to some extent, well-mixed, which lead to inaccuracies when trying to determine in which conditions spatial patterns form in vegetation systems. The spatial moment model provides a middle ground, balancing computational feasibility with the ability to incorporate spatial structure explicitly. This approach retains key insights from IBMs regarding spatial patterns and spatially explicit interactions while remaining computationally tractable for larger-scale analyses.